# Stepwise Reasoning Error Disruption Attack of LLMs

### v1

@@ Abstract
Large language models (LLMs) have made remarkable strides in complex reasoning tasks, but their safety and robustness in reasoning processes remain underexplored. Existing attacks on LLM reasoning are constrained by specific settings or lack of imperceptibility, limiting their feasibility and generalizability. To address these challenges, we propose the Stepwise rEasoning Error Disruption (SEED) attack, which subtly injects errors into prior reasoning steps to mislead the model into producing incorrect subsequent reasoning and final answers. Unlike previous methods, SEED is compatible with zero-shot and few-shot settings, maintains the natural reasoning flow, and ensures covert execution without modifying the instruction. Extensive experiments on four datasets across four different models demonstrate SEED’s effectiveness, revealing the vulnerabilities of LLMs to disruptions in reasoning processes. These findings underscore the need for greater attention to the robustness of LLM reasoning to ensure safety in practical applications.
@@
Jingyu Peng‡§111 Both authors contributed equally to this research., Maolin Wang§111 Both authors contributed equally to this research., Xiangyu Zhao§ 222 Corresponding authors. , Kai Zhang, Wanyu Wang§, Pengyue Jia§, Qidong Liu§, Ruocheng Guo, Qi Liu 222 Corresponding authors. University of Science and Technology of China, § City University of Hong Kong, ByteDance Research jypeng28@mail.ustc.edu.cn morin.wang@my.cityu.edu.hk
@@ 1 Introduction
Previous work has exposed significant LLM vulnerabilities in simple tasks such as classification and generation Wang et al. (2024a); Zhao et al. (2023); Xu et al. . However, their susceptibility to attacks during the complex reasoning processes - where the stakes are often higher and the consequences are more severe in some critical area - remains largely unexplored. Recent advances in reasoning methods have introduced effective mechanisms, enabling models to iteratively build upon prior steps: from basic chain-of-thought approaches Wei et al. (2022); Kojima et al. (2022) to advanced reflection mechanisms  Madaan et al. (2024); Zhao et al. (2024) and multi-step reasoning Wu et al. (2024); Wang et al. (2024b). This critical dependence on step-wise reasoning introduces a new type of vulnerability in LLMs, where manipulation of initial reasoning steps can propagate errors, causing cascading failures throughout the reasoning chain.
@@
Exploiting such vulnerability in LLMs introduces two fundamental challenges: feasibility and imperceptibility. Technically, unlike traditional adversarial attack methods, which often leverage internal information of models such as gradients and logits, state-of-the-art LLMs are now primarily deployed as proprietary APIs Achiam et al. (2023); Team et al. (2023). Therefore, only prompt-based attacks are feasible, where adversaries have to operate through input manipulation. While existing attempts to compromise LLM reasoning Xu et al. (2024); Xiang et al. (2024) have demonstrated success in specific scenarios, they still face severe limitations in practice. A key challenge in attack design is crafting imperceptible attacks. While obvious manipulations like changing final answers or inserting irrelevant steps are easily detected, modifying the reasoning output while maintaining narrative coherence is rather challenging. Existing approaches struggle to achieve this balance between attack effectiveness and stealth in complex reasoning tasks.
@@
Among the most relevant approaches, Xiang et al. (2024) employs misleading demonstrations to induce errors in LLMs. However, these methods are limited to in-context learning scenarios, requiring demonstrations as input, which limits their generalizability to the zero-shot settings. Furthermore, their strategy introduces an additional step that modifies the final answer, making it quite easy to identify. Another related approach, the preemptive answer “attack” Xu et al. (2024) alters the reasoning paradigm of the model by producing conclusions before deriving reasoning steps. Despite its novelty, this approach often generates easily identifiable outputs, reducing its imperceptibility and effectiveness in practice. These limitations are further evidenced by our experimental results in Section 3.2.
@@
  • We define the task of disrupting the step-by-step reasoning process of LLMs and introduce SEED, a versatile and effective attack method that is both highly efficient in execution and challenging in detection.
@@ 2 Method
Figure 2: An example of the SEED-S attack on a math problem. The red font highlights the generated misleading content where subtle calculation errors are introduced while maintaining the reasoning coherence.
@@ 2.1 Problem Formulation
We first present a formal definition of a step-by-step reasoning task of LLMs as shown in Fig 1. For a given problem $p$, we define the query to the LLM, denoted as $q$, as follows:
@@
Figure 3: An example of the SEED-P attack on a math problem. The red font highlights the generated misleading content, while the gray font highlights the optional part for problems with answer choices.
@@ 2.2 Overview of Stepwise Reasoning Error Disruption Attack
with the constraint that $T_{att}+T^{\prime}=T$. In practice, as the number of reasoning steps $T$ varies, we introduce $\sigma=\frac{T_{att}}{T}$ as a hyperparameter to control the $T_{att}$. To generate $R_{att}$ that both closely resembles $R[:T_{att}]$ and effectively misleads the LLM into providing an incorrect answer, we developed two LLM-based implementations. In the following two subsections, we propose two implementations of SEED attack: SEED-S (Step Modification) and SEED-P (Problem Modification). SEED-S directly modifies the final step of the reasoning process, while SEED-P generates a modified problem that leads to the desired incorrect answer.
@@ 2.3 SEED-S: SEED Attack by Step Modification
However, this naive implementation has some drawbacks. Firstly, it has been observed that LLMs tend to pay more attention to the head and tail parts of the input. As a result, LLMs are more likely to notice inconsistencies between the last two steps. Secondly, the final result $a^{\prime}$ is easily noticeable, for example, if there is a significant discrepancy in magnitude compared to $a$, or if the final answer in a multiple-choice question does not appear among the available options. In such cases, the LLM is likely to recognize and correct the error.
@@
2.5 Properties Analysis
Both implementations demonstrate unique characteristics in achieving adversarial objectives. SEED-S, by directly modifying the final reasoning step, alters only a small portion of the input, resulting in weaker attack capabilities but making it less detectable. In contrast, SEED-P demonstrates stronger attack effectiveness by generating a fully modified problem with coherent reasoning chains, showcasing broader applicability across various problem types, including multiple-choice questions and open-ended reasoning tasks.
@@ 3.1 Experimental Setup
Dataset. Building on prior studies targeting reasoning processes in LLMs Xu et al. (2024); Xiang et al. (2024), we evaluate our method using four datasets that encompass diverse and challenging reasoning tasks presented in two formats. Specifically, MATH Hendrycks et al. and GSM8K Cobbe et al. (2021) focus on arithmetic reasoning with open-ended formats, while MATHQA Amini et al. (2019) presents math problems in a multiple-choice format. CSQA Talmor et al. (2019), on the other hand, features multiple-choice commonsense reasoning tasks. As for the budget constraints, we follow the approach of Xiang et al. (2024), randomly sampling 500 questions from each dataset for our experiments. Further details about datasets are provided in Appendix B.
@@
LLMs. We evaluate four cutting-edge LLMs, encompassing both open-source and proprietary models: Llama3-8B Dubey et al. (2024), Qwen-2.5-7B Hui et al. (2024), Mistral-v0.3-7B Jiang et al. (2023), and GPT4-o Achiam et al. (2023). These models are chosen for their state-of-the-art performance and strong capabilities in solving complex reasoning tasks, providing a comprehensive benchmark to evaluate the effectiveness and versatility of our proposed attack methodology.
@@
Settings. To assess the generalizability of SEED attack, we test its performance in both zero-shot and few-shot settings, following the traditional prompt-based Chain-of-Thought (CoT) paradigm Wei et al. (2022); Kojima et al. (2022). In the main experiments, we set $\sigma$ to 0.6, and the impact of varying $\sigma$ is explored in Section 3.3. Our experiments ’ technical specifications and implementation details are available in Appendix D.
@@
Metrics. We assess the performance using three key metrics: accuracy (ACC), attack success rate (ASR) and detection rate. ACC measures the percentage of problems solved correctly by the model. ASR quantifies the proportion of originally correct answers that are rendered incorrect due to the attack, serving as a direct indicator of the attack’s effectiveness in disrupting the model’s reasoning capabilities. The Detection rate measures the proportion of solutions identified as originating from attacked input prompts. Further information on the metrics is available in the Appendix A.
Baselines. To the best of our knowledge, the only comparable methods are UPA and MPA proposed by Xu et al. (2024). Both approaches instruct the LLM to provide the answer first, followed by the reasoning steps. MPA goes a step further by introducing an LLM-generated false answer to mislead the model into producing corresponding incorrect reasonings. Although BadChain Xiang et al. (2024) makes a notable contribution by achieving an ASR close to 100% across all datasets, its applicability is restricted to the few-shot setting. Furthermore, as shown in Table 1, its detection ratio approaches 100%, as it simply modifies the final answer. Due to this limitation, BadChain is excluded from subsequent discussions.
@@
SEED 0.453 0.311 0.409 0.553
@@
3.2 Overall Performance
Evaluation on Covert Detection As outlined in Section 1, we hypothesize that modifying $I_{solve}$ will lead to LLM outputs that are more easily identifiable, thereby diminishing the covert nature of the attack. To test this hypothesis, we evaluated the detection rates of SEED-generated solutions against baseline methods using GPT4-o. Specifically, GPT4-o was employed to classify whether the input prompt had been attacked by analyzing the corresponding output solutions. Table 1 presents detection rates on the MATH and GSM8K datasets, demonstrating that UPA and MPA fail to satisfy the covert requirements. In contrast, both SEED-S and SEED-P exhibit substantial improvements in detection rates, particularly with GPT4-o, where the average improvement exceeds 90%. As a result, GPT4-o is substantially less likely to detect the outputs of attacked LLMs under SEED attack, validating that SEED achieves a higher level of stealth by maintaining a natural reasoning flow without apparent manipulations.
Performance Comparison in Baseline Settings To ensure a fair evaluation of effectiveness, we adapted the SEED attack to match the same settings as UPA and MPA, incorporating instructions for the LLM. As shown in Table 2, SEED attack achieves improved attack performance in most cases, compared to UPA and MPA. The performance gap on CSQA is especially evident. On GPT-4o, SEED achieved an ASR more than 2x that of the baseline. This is due to the inclusion of additional reasoning steps in SEED attack that further enhance attack performance compared to UPA and MPA in most cases, indicating that SEED attack is compatible with UPA or MPA. This improvement is attributed to the SEED attack’s ability to introduce additional reasoning steps. Furthermore, these results demonstrate that SEED attack is not only a standalone approach but also compatible with methods like UPA and MPA, potentially offering a hybrid strategy for further enhancing attack performance.
Table 3: Performance comparison of the two SEED attack variations, evaluated using ACC (Accuracy) and ASR (Attack Success Rate). SEED-S and SEED-P denote SEED attack implemented through step modification and problem modification, respectively. Lower ACC and higher ASR indicate a greater impact of SEED attack. Method N represents the raw performance without any attack. Lowest ACC and highest ASR are highlighted.
@@ 3.2 Overall Performance
Effectiveness Evaluation We evaluated the effectiveness of the two implementations on all datasets and all models. As shown in Table 3, although the results vary significantly across different datasets and models, all LLMs demonstrate vulnerability under the proposed SEED attack, substantially reducing the original ACC across all tasks under both Zero-shot and Few-shot settings.
SEED-S occasionally fails, as demonstrated in CSQA and MATHQA, where the ASR is only 0.069 and 0.064 under few-shot setting, respectively. In contrast, SEED-P consistently outperforms SEED-S across all tasks. Specifically, in the CSQA and MATHQA multiple-choice datasets, SEED-S shows a low attack success rate, while SEED-P, with its more effective design, substantially boosts the ASR and reduces ACC.
Comparing the performance across different models, we find that Qwen and GPT-4o show greater robustness to the SEED attack than other models, particularly GPT-4o on MATH and GSM8K, and Qwen on CSQA and MATHQA, with ASR all under 0.4. Additionally, these models exhibit relatively higher original accuracy on the corresponding datasets. Based on these observations, we infer that a model’s performance and robustness on a task are positively correlated.
To validate this inference, we further conduct SEED-P on questions that are answered correctly and incorrectly independently. The results shown in Table 4 reveal a significant ASR gap between the two groups. The largest gap is observed in Llama-3 under the few-shot setting, achieving an ASR of 0.417. This suggests that LLMs are more robust on questions that they are capable of answering correctly, which aligns with our intuition. We also conducted experiments to investigate the transferability of SEED-P; due to space constraints, the results are presented in Appendix E.

### v3

@@ Abstract
Large language models (LLMs) have made remarkable strides in complex reasoning tasks, but their safety and robustness in reasoning processes remain unexplored, particularly in third-party platforms that facilitate user interactions via APIs. Existing attacks on LLM reasoning are constrained by specific settings or lack of imperceptibility, limiting their feasibility and generalizability. To address these challenges, we propose the Stepwise rEasoning Error Disruption (SEED) attack, which subtly injects errors into prior reasoning steps to mislead the model into producing incorrect subsequent reasoning and final answers. Unlike previous methods, SEED is compatible with zero-shot and few-shot settings, maintains the natural reasoning flow, and ensures covert execution without modifying the instruction. Extensive experiments on four datasets across four different models demonstrate SEED’s effectiveness, revealing the vulnerabilities of LLMs to disruptions in reasoning processes. These findings underscore the need for greater attention to the robustness of LLM reasoning to ensure safety in practical applications. Our code is available at: https://anonymous.4open.science/r/Reasoning-Attack-4663/
@@
Jingyu Peng‡§11footnotemark: 1, Maolin Wang§11footnotemark: 1, Xiangyu Zhao§ 22footnotemark: 2 , Kai Zhang, Wanyu Wang§, Pengyue Jia§, Qidong Liu§, Ruocheng Guo, Qi Liu 22footnotemark: 2 University of Science and Technology of China, § City University of Hong Kong, Independent Researcher jypeng28@mail.ustc.edu.cn morin.wang@my.cityu.edu.hk
@@ 1 Introduction
In practice, LLMs are increasingly deployed through third-party platforms that mediate user interactions via APIs, where users do not directly access the models. This setup introduces a security risk: malicious providers could manipulate reasoning or outputs—even if model outputs seem normal at first glance, resulting in incorrect reasoning and conclusions. In this work, we investigate this specific risk by focusing on how these platforms might compromise model integrity by input manipulation.
@@
Previous work has exposed significant LLM vulnerabilities in simple tasks such as classification and generation Wang et al. (2024); Zhao et al. (2023); Xu et al. . However, their susceptibility to attacks during the complex reasoning processes—where the stakes are often higher and the consequences are more severe in some critical areas—remains largely unexplored.
Recent advances in long reasoning methods require LLMs to iteratively build upon prior steps, facilitating reflection Madaan et al. (2024); Zhao et al. (2024) or tree search Guan et al. (2025); Zhang et al. (2024) for subsequent reasoning steps. This critical dependence on step-wise reasoning introduces a new type of vulnerability in LLMs, where manipulation of initial reasoning steps can propagate errors, causing cascading failures throughout the reasoning chain.
@@
Exploiting such vulnerability in LLMs introduces two fundamental challenges: feasibility and imperceptibility. Technically, unlike traditional adversarial attack methods, which often leverage internal information of target models such as gradients and logits, state-of-the-art LLMs are now primarily deployed as proprietary APIs Achiam et al. (2023); Team et al. (2023). Therefore, only prompt-based attacks are feasible, where adversaries have to operate through input manipulation. While existing attempts to compromise LLM reasoning Xu et al. (2024); Xiang et al. (2024) have demonstrated success in specific scenarios, they still face severe limitations in practice. A key challenge in attack design is to create attacks that are imperceptible to users. While obvious manipulations, such as altering final answers or inserting irrelevant steps, are easily detected by users, modifying the reasoning process while preserving narrative coherence is far more difficult. Existing methods often struggle to balance attack effectiveness with stealth, especially in the context of complex reasoning tasks.
@@
Among the most relevant approaches, Xiang et al. (2024) employs misleading demonstrations to induce errors in LLMs. However, these methods are limited to in-context learning scenarios, requiring demonstrations as input, which limits their generalizability to the zero-shot settings. Furthermore, their strategy introduces an additional step that modifies the final answer, making it quite easy to identify by users. Another related approach, the preemptive answer “attack” Xu et al. (2024) alters the reasoning paradigm of the model by producing conclusions before deriving reasoning steps. Despite its novelty, this approach often generates easily identifiable outputs, reducing its imperceptibility and effectiveness in practice. These limitations are further evidenced by our experimental results in Section 3.2.
@@
  • We define the task of disrupting the step-by-step reasoning process of LLMs and introduce SEED, a versatile and effective attack method that is both highly efficient in execution and challenging in detection by users.
@@ 2 Method
Figure 2: An example of SEED-S/P attacks on a math problem. The red font highlights misleading content, where subtle calculation errors are introduced while preserving reasoning coherence. SEED-S/P first generates the initial reasoning steps containing errors, after which the target LLM produces the subsequent steps. These components are seamlessly integrated to present a complete and coherent chain-of-thought reasoning process to the user.
@@ 2.1 Problem Formulation
We first present a formal definition of a step-by-step reasoning task of LLMs as shown in Figure 1. For a given problem $p$, we define the query to the LLM, denoted as $q$, as follows:
@@ 2.2 Overview of Stepwise Reasoning Error Disruption Attack
with the constraint that $T_{att}+T^{\prime}=T$. In practice, as the number of reasoning steps $T$ varies, we introduce $\sigma=\frac{T_{att}}{T}$ as a hyperparameter to control the $T_{att}$. To generate $R_{att}$ that both closely resembles $R[:T_{att}]$ and effectively misleads the LLM into providing an incorrect answer, we developed two LLM-based implementations.
In the next two subsections, we introduce two implementations of the SEED attack: SEED-S (Step Modification) and SEED-P (Problem Modification). SEED-S directly alters the final step of the reasoning process, whereas SEED-P modifies the problem itself to produce the desired incorrect answer.
@@ 2.3 SEED-S: SEED Attack by Step Modification
However, this naive implementation has a significant limitation in terms of attack effectiveness. First, it has been observed that LLMs tend to focus more on the beginning and end of the input. As a result, they are more likely to detect inconsistencies in the final steps. Additionally, altering just a single reasoning step is often insufficient to convincingly mislead the target LLM.
@@
It’s worth noting that although SEED-P requires $LLM_{assist}$ to initially answer the question, the accuracy of the answer has limited impact on the SEED-P’s performance. For short-answer questions, SEED-P remains effective regardless of the initial answer’s accuracy, successfully introducing faulty reasoning steps across various model performance levels. For multiple-choice questions, let the accuracy of the LLM’s responses be denoted as $p$, with a total of $K$ options for each question. While we acknowledge the theoretical constraint that the attack failure probability is $(1-p)\cdot\frac{1}{K-1}$, its effect on the model’s overall attack ability is still relatively minimal.
@@ 3.1 Experimental Setup
Dataset. Building on prior studies targeting reasoning processes in LLMs Xu et al. (2024); Xiang et al. (2024), we evaluate our method using four datasets that encompass diverse and challenging reasoning tasks presented in two formats. Specifically, MATH Hendrycks et al. and GSM8K Cobbe et al. (2021) focus on arithmetic reasoning with open-ended formats, while MATHQA Amini et al. (2019) presents math problems in a multiple-choice format. CSQA Talmor et al. (2019), on the other hand, features multiple-choice commonsense reasoning tasks. As for the budget constraints, we follow the approach of Xiang et al. (2024), randomly sampling 500 questions from each dataset for our experiments. Further details about datasets are provided in Appendix A.
@@
Backbone LLMs. We evaluate four cutting-edge LLMs, encompassing both open-source and proprietary models: Llama3-8B Dubey et al. (2024), Qwen-2.5-7B Hui et al. (2024), Mistral-v0.3-7B Jiang et al. (2023), and GPT-4o Achiam et al. (2023). These models are chosen for their state-of-the-art performance and strong capabilities in solving complex reasoning tasks, providing a comprehensive benchmark to evaluate the effectiveness and versatility of our proposed attack methodology.
@@
Settings. To assess the generalizability of SEED attack, we test its performance in both zero-shot and few-shot settings, following the traditional prompt-based Chain-of-Thought (CoT) paradigm Wei et al. (2022); Kojima et al. (2022). In the main experiments, we set $\sigma$ to 0.6, and the impact of varying $\sigma$ is explored in Section 3.3. Our experiments’ technical specifications and implementation details are available in Appendix B.
@@
Metrics. We assess the performance using four key metrics: accuracy (ACC), attack success rate (ASR), modification success rate (MSR) and detection rate. ACC measures the percentage of problems solved correctly by the model. ASR quantifies the proportion of originally correct answers that are rendered incorrect due to the attack, serving as a direct indicator of the attack’s effectiveness in disrupting the model’s reasoning capabilities. MSR quantifies the proportion of problems that are altered by the attack. The detection rate measures the proportion of solutions identified as originating from attacked input prompts. Further information on the metrics is available in the Appendix C.
Baselines. To our knowledge, UPA and MPA, introduced by Xu et al. (2024), along with BadChain Xiang et al. (2024), are the only methods targeting attacks on LLM reasoning. UPA and MPA prompt the LLM to generate an answer before the reasoning steps, with MPA further introducing a false answer to mislead reasoning. While BadChain achieves an ASR close to 100% across all datasets, its effectiveness is limited to the few-shot setting. Moreover, as Table 1 shows, its detection ratio nears 100% since it only modifies the final answer, warranting its exclusion from further discussion. Additionally, we find that the “Adding Mistake” method in Lanham et al. (2023) shares similarities with SEED-S, in that it introduces misleading reasoning steps. However, the “Adding Mistake” approach primarily focuses on examining whether CoT reasoning is post-hoc, rather than attack the reasoning of LLMs. Since the task of “Adding Mistake” differs from our single-round reasoning task, we concentrate solely on comparing the effectiveness of the attack.
@@
SEED-P 0.453 0.311 0.409 0.553
3.2 Overall Performance
Evaluation on Covert Detection
As outlined in Section 1, we hypothesize that modifying $I_{solve}$ will lead to LLM outputs that are more easily identifiable by human, thereby diminishing the covert nature of the attack. To test this hypothesis, we evaluated the detection rates of SEED-generated solutions against baseline methods using GPT-4o to simulate human detection following the LLM-as-a-judge paradigm (Gu et al., 2024). Specifically, GPT-4o was employed to classify whether the input prompt had been attacked by analyzing the corresponding output solutions. Table 1 presents detection rates on the MATH and GSM8K datasets, demonstrating that UPA and MPA fail to satisfy the covert requirements. In contrast, both SEED-S and SEED-P exhibit substantial improvements in detection rates, particularly with GPT-4o, where the average improvement exceeds 90%. This validates that SEED achieves a higher level of stealth by preserving a natural reasoning flow without apparent manipulations.
To closely simulate real-world scenarios, we conducted a rigorous human evaluation study involving 20 PhD students in computer science. Using the MATH and CSQA datasets, as detailed in Appendix D, their assessments closely aligned with our findings from GPT-4o.
@@ 3.2 Overall Performance
Performance Comparison in Baseline Settings To ensure a fair evaluation of effectiveness, we adapted the SEED-P attack to match the same settings as UPA and MPA, incorporating instructions for the LLM. As shown in Table 2, SEED-P attack achieves improved attack performance in most cases, compared to UPA and MPA. The performance gap on CSQA is especially evident. On GPT-4o, SEED-P achieved an ASR more than 2x that of the baseline. This is due to the inclusion of additional reasoning steps in SEED-P attack that further enhance attack performance compared to UPA and MPA in most cases, indicating that SEED-P attack is compatible with UPA or MPA. This improvement is attributed to the SEED-P attack’s ability to introduce additional reasoning steps. Furthermore, these results demonstrate that SEED-P attack is not only a standalone approach but also compatible with other methods like UPA and MPA, potentially offering a hybrid strategy to further enhance attack performance.
Table 3: Performance comparison of the two SEED attack variations and “Adding Mistake” in Lanham et al. (2023), evaluated using ACC (Accuracy) and ASR (Attack Success Rate). SEED-S and SEED-P denote SEED attack implemented through step modification and problem modification, respectively. Lower ACC and higher ASR indicate a greater impact of SEED attack. Method N represents the raw performance without any attack. Lowest ACC and highest ASR are highlighted.
@@
Effectiveness Evaluation
We evaluated the effectiveness of SEED implementations and the “Adding Mistake” method across various datasets and models. As shown in Table 3, although results vary, all LLMs are vulnerable to the SEED attack, significantly reducing ACC in both zero-shot and few-shot settings. SEED-S and "Adding Mistake" perform similarly, but SEED-S generally has higher attack success rates in most cases, likely due to the summarization step in “Adding Mistake” that may alert the model to inconsistencies. SEED-S occasionally fails due to its limited ability, as seen in CSQA and MATHQA, with ASR of 0.069 and 0.064 in few-shot settings. However, SEED-P consistently outperforms SEED-S across all tasks, particularly in the CSQA and MATHQA datasets, where SEED-P greatly increases ASR and reduces ACC. This improvement is due to $LLM_{assist}$’s ability to adapt to different problems and modify key elements affecting outcomes, as shown in Appendix E.
Comparing the performance across different models, we find that Qwen and GPT-4o are more robust to the SEED attack than other models, particularly GPT-4o on MATH and GSM8K, and Qwen on CSQA and MATHQA, with ASR values all under 0.4. Additionally, these models exhibit relatively higher original accuracy on the corresponding datasets, suggesting a positive correlation between a model’s performance and robustness on a task. To validate this, we applied SEED-P separately to questions the LLM answers correctly and incorrectly, then evaluated the MSR independently. Results in Table 4 show a significant MSR gap between the two groups, with the largest gap in Llama-3 under the few-shot setting, reaching an MSR of 0.417. This indicates that LLMs are more robust on questions they answer correctly, aligning with our inference. Furthermore, the transferability evaluation presented in Appendix F confirms that more powerful LLMs can achieve both a high ASR as the assistant LLM and strong robustness as the target LLM.
In Appendix G, we evaluate self-review prompts under zero-shot settings, finding only modest improvements with ASR decreasing by no more than 10%. This suggests that simple prompt-based defenses need further refinement to counter SEED attacks. We also validated the effectiveness of prepending a wrong answer and 2-stage reasoning step generation by conducting an ablation study (see Appendix H).